OpenStudy (anonymous):

The height of the triangle is 4 cm. more than twice the4 length of the base.The area of the triangle is 35 cm.^2. Find the height of the triangle

OpenStudy (anonymous):

This is how I would approach it. You know the height=4cm+2b, AofTri= 35cm^2, and that area of a triangle is 1/2b*h. So we have, 35=1/2(4+2b)b. 35=(2+b)b=2b+b^2. Write it as a quadratic equation. 0=b^2+2b-35. Factor and get (x+7)(x-5). Your zeros are -7 and 5. You know you can't have a negative base value so your base in this case is 5. Plug and chug to see if your answer is right!